OpenStudy (anonymous):

According to the Rational Root Theorem, which function has the same set of potential rational roots as the function g(x) = 3x^5 – 2x^4 + 9x^3 – x^2 + 12?

OpenStudy (anonymous):

f(x) = 3x^5 – 2x^4 – 9x^3 + x^2 – 12 f(x) = 3x^6 – 2x^5 + 9x^4 – x^3 + 12x f(x) = 12x^5 – 2x^4 + 9x^3 – x^2 + 3 f(x) = 12x^5 – 8x^4 + 36x^3 – 4x^2 + 48

OpenStudy (misty1212):

pick the one with the same leading coefficient and constant (ignoring the signs)